find the maximum and minimum points of the function x3-x2

1
Expert's answer
2021-09-06T16:31:58-0400
"y=x^3-x^2"

Domain: "(-\infin, \infin)"

Find the first derivative


"y'=(x^3-x^2)'=3x^2-2x"

Find the critical point(s)


"y'=0=>3x^2-2x=0"

"x(3x-2)=0"

"x_1=0, x_2=\dfrac{2}{3}"

Critical numbers: "0, \dfrac{2}{3}."

First Derivative Test

If "x<0," then "y'>0, y" increases.

If "0<x<\dfrac{2}{3}," then "y'<0, y" decreases.

If "x>\dfrac{2}{3}," then "y'>0, y" increases.


"y(0)=(0)^3-(0)^2=0"

"y(0)=(\dfrac{2}{3})^3-(\dfrac{2}{3})^2=-\dfrac{4}{27}"

The function "y=x^3-x^2" has a local maximum with value of at "x=0,"

"Point(0, 0)."

The function "y=x^3-x^2" has a local minimum with value of "\dfrac{2}{3}" at "x=-\dfrac{4}{27},"

"Point(-\dfrac{2}{3}, -\dfrac{4}{27})."


Since the function "y=x^3-x^2" is defined on "\R," the function "y" has neither absolute maximum nor absolute minimum.